OpenStudy (redheadangel):

A cylinder’s radius is reduced to 2/5 its original size and the height is quadrupled. How has the volume of the cylinder changed?

OpenStudy (redheadangel):

I think I know the answer but I just wanted a second opinion

OpenStudy (redheadangel):

@Peaches15

OpenStudy (redheadangel):

@CausticSyndicalist

OpenStudy (anonymous):

well if you let radius = 1 and height = h the volume of the cylinder is \[V = \pi \times1^1\times h = \pi h\] so if the radius becomes 2/5 and height becomes 4h then \[V = \pi \times (\frac{2}{5})^2 \times 4h\] or \[V=\frac{16}{25} \pi h\]

OpenStudy (redheadangel):

Thank you soooo much!!